OpenStudy (yessenia27):

An ice sculpture is melting from the heat. Assuming its height changes by -5/32 meters every hour, what will be the change in height after 4 hours?

OpenStudy (mww):

If the rate of change is constant, then rate of change = change in quantity / time So the change in quantity = rate of change x time

jhonyy9 (jhonyy9):

use the rule of 3 simple 1 h ---------------- -5/32 m 4 h --------------- x ------------------------- x = 4(-5/32)/1 = 4*(-5/32) = -5/8 m x = -5/8 m